For the electrical circuit below, let R1=30 Ohms, R2=15 Ohms, R3=10 Ohms, R4=20 Ohms, C=0.1 F and Vin=10V and complete the following.For the electrical circuit below, let R1 = 30 12, R2 = 152, R3 = 102, R1 = 20 2, C =0.1 F and Vin = 10V and complete the foll
Answer #1

Given, R,= 3on Ra=15n Rz=lon c=o.lt Vin lov R4 =2on. Circuit Rz-lon 0 R,=3an R2 = 150 + 3 Rudon vout =lov recht Apply Nodal aact) 3 dvect) -+ avect vCt) - 10 30 tie Ct) + vet) 30 vct tic Ct 3 cdy(t) dt (c=01F) te dvect) اليا I dvc (t) = 1 dt =) VCA dToansto tor & Apply Apply lap ducat? - Sve(s) 5 35 VCCs) (for Resocuits duelt) + 1 vect) = 5 - Sve(s) + 1 ve(s) = 1/2 ve(s) €we got Vc Ct) L = 3.16 3 [i- ést at to, vect) = 5(1-1)=0 tar, vect) = 5(1-03 time constant = 5 [l-el too, Vect) = 5 [1-6] = 5[=0.If Capacitor Curdent, ic Ct)= c dvect) dt we got welt) = 5 (-est ic Ct) = 1 x5 dci-e-Ist] 10 at It 11 2 iCt)= icct)- 10 - loe & t. C = -56-1 ezt time taken to deach its final value is, (v (t)) T= time constant = 5x3 = 3sec 15 sec from circuAt t=0, vout (t) = 10-25 At t-T, vout (A) - le 50 10-5e .6.9343 At tno vout Ct) Esteady state rout(t) value) 6.343 con اد T-